OpenStudy (alyssia):

. x = y + 6, y = -2 - x

OpenStudy (anonymous):

x=2 and y=-4

OpenStudy (ash2326):

We have \[x=y+6\] and \[y=-2-x\] from second equation we have y=-2-x, let's substitute this in first equation, we get \[x=(-2-x)+6\] or \[x=-2-x+6\] add both sides by x \[2x=-2+6\] or \[2x=4\] divide both sides by 2, we get \[x=2\] now we have \[y=-2-x\] let's substitute x=2, we get \[y=-2-2=-4\] so we egt \[x=2 , y=-4\]